OpenStudy (jazzyfa30):

Pre cal help f(x)=sqrrt 3x+4, G(x)=sqrrt 16x-9 find f*g

OpenStudy (jazzyfa30):

the answer I got was (-3/4)

OpenStudy (jazzyfa30):

@amistre64

OpenStudy (solomonzelman):

\[\sqrt{3x+4~} ~\times~ \sqrt{16x-9~}\]\[\sqrt{~~(3x+4)~(16x-9)~~}\]\[\sqrt{~~48x^2+21x-36~~}\] that's what I got.